OpenStudy (anonymous):

use the intercept to graph the equation x+3y=6

myininaya (myininaya):

write in y=mx+b form so subtract x on both sides: 3y=-x+6 now divide both sides by 3: y=-x/3+2 so b=2 m=-1/3 

myininaya (myininaya):

see how i graph the y-intercept (0,2) then i counted down once and over three times to the right using m